Grade 12th passGeneral Physics

the velocity v and displacement r of a body are related as v2=kr, where k is constant. what will be trhe velocity after 1 second/ (given that displacement is zero at t=0)

adyasha
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
 v2=kr if compared with
v^2 =  u^2 + 2as
then u = 0
a = k/2


Velocity after 1 second = >>
v = u + at
v = 0  + k/2t = k/2*(1) = 0.5k m/s